OpenStudy (anonymous):

I've had problems understanding situations involving the mixture of substances that are already in a ratio. I can't work with two or more ratios. Here's an example sum that fails me - "There are three containers of equal capacity and all are completely filled with acid and water mixtures in different ratios. The ratio of acid to water in the first container is 2:3, in the second container is 3:7 and in the third container it is 4:11. If the mixtures of all the three containers are mixed together, what is the ratio of acid to water in it?"

OpenStudy (anonymous):

since you are not told the amount in each, a nice gimmick to use is to pick a number, compute with that number, and then since the number is unimportant, the answer will be correct

OpenStudy (anonymous):

i would pick the least common multiple of 3, 7 and 11, which is 231

OpenStudy (anonymous):

then \(\frac{2}{3}\times 231 =154, \frac{3}{7}\times 231=99, \frac{4}{11}\times 231=84\)

OpenStudy (anonymous):

But the answer is 29/61

OpenStudy (anonymous):

I'll try to explain my method

OpenStudy (anonymous):

the 3 ratios given are 2:3, 3:7 and 4:11 Based on each ratio, add them up to find the total composition of each ratio That would give 5, 10 and 15 Scale these numbers to a common multiple. 30 is good. so, multiply the first ratio by 6, the second by 3 and the thord by 2

OpenStudy (anonymous):

that would make the three ratios as 12:18, 9:21 AND 8:22

OpenStudy (anonymous):

add the components of the acid and water from the three ratios

OpenStudy (anonymous):

you get 29:61

OpenStudy (anonymous):

oh yeah i am way off sorry

OpenStudy (anonymous):

it is five parts, ten parts and 15 parts, not what i wrote

OpenStudy (anonymous):

so make them all 30 parts as said above

OpenStudy (anonymous):

Irkiz thanks a lot